2. During a sale at a men’s store, 16 white sweaters, 3 red sweaters, 9 blue sweaters, and 7 yellow sweaters were purchased. If a customer is selected at random, find the probability that he bought,


a. A blue sweater. (3 marks)


b. A yellow or a white sweater. (3 marks)


c. A red, blue or a yellow sweater. (3 marks)


d. A sweater that was not white (3 marks)




1
Expert's answer
2022-03-29T07:40:13-0400

"a:p=\\frac{9}{16+3+9+7}=0.257143\\\\b:p=\\frac{7+16}{16+3+9+7}=0.657143\\\\c:p=\\left[ \\begin{array}{c}\tnot\\\\\twhite\\\\\\end{array} \\right] =1-\\frac{16}{16+3+9+7}=0.542857\\\\d:p=\\left[ the\\,\\,same\\,\\,as\\,\\,c \\right] =0.542857\\\\"